# LogPTR: Variable-Aware Log Parsing with Pointer Network

### 11 Jan 2024 v1

@@ 4.1 Effectiveness on General Log Parsing
Android
0.916
0.943
Apache
0.993
1
BGL
0.896
0.959
Hadoop
0.968
0.968
HDFS
0.965
0.998
HealthApp
0.988
0.971
HPC
0.99
0.992
Linux
0.959
0.961
Mac
0.862
0.866
OpenSSH
0.976
0.994
OpenStack
0.932
0.998
Proxifier
1
1
Spark
0.991
0.997
Thunderbird
0.878
0.925
Windows
0.99
0.983
Zookeeper
0.981
0.989
Average
0.955
0.972
4.2 Effectiveness on Variable-aware Log Parsing
Apart from identifying static and dynamic parts in logs (i.e., general log parsing), LogPTR can also identify the categories of variables (i.e., variable-aware log parsing). In this RQ, we compare the accuracy of LogPTR on variable-aware log parsing with VALB [15], which is the only baseline that can identify the categories of variables. Table 2 presents the accuracy of LogPTR and VALB on variable-aware log parsing.
From the results, we can see that LogPTR achieves a high parsing accuracy ranging from 0.866 in Mac to 1 in Apache, which is also close to the parsing accuracy as discussed in RQ1. The average parsing accuracy of LogPTR is 0.972, which is higher than the baseline (i.e., 0.955). Moreover, LogPTR achieves the best results on 14 out of 16 datasets. As for the robustness, we calculate the standard deviation (STD) of parsing accuracy, and LogPTR achieves a smaller STD than the baseline (0.036 versus 0.045). Overall, the experimental results confirm that apart from general log parsing, LogPTR can also efficiently identify the categories of variables in logs to perform variable-aware log parsing.
@@ 5 CONCLUSION
This paper proposes LogPTR, the first end-to-end variable-aware log parser that can extract static and dynamic parts in logs, and further identify categories of variables. The key of LogPTR is using pointer network to copy words from the log message. Experimental results on 16 public log datasets show that LogPTR outperforms the state-of-the-art log parsers both on general log parsing and variable-aware log parsing.

### 13 Mar 2026 v2

@@ 4.2 Effectiveness on Variable-aware Log Parsing
Android 91.6 94.3
Apache 99.3 100
BGL 89.6 95.9
Hadoop 96.8 96.8
HDFS 96.5 99.8
HealthApp 98.8 97.1
HPC 99.0 99.2
Linux 95.9 96.1
Mac 86.2 86.6
OpenSSH 97.6 99.4
OpenStack 93.2 99.8
Proxifier 100 100
Spark 99.1 99.7
Thunderbird 87.8 92.5
Windows 99.0 98.3
Zookeeper 98.1 98.9
Average 95.5 97.2
5 RELATED WORK
Existing Log parsers can be categorized into syntax-based and semantic-based log parsers. Syntax-based log parsers leverage manually designed features to extract log templates, which have been widely explored in the past. Specifically, syntax-based log parsers can be further subdivided into three categories. (1) Frequency-based parsers: These log parsers [23, 19, 22, 2] utilize frequent patterns of token position or n-gram information to distinguish the templates and variables in log messages. (2) Clustering-based parsers: These log parsers [5, 21, 20, 6] compute similarities between log messages to cluster them into different groups and then extract the constant parts of log messages. (3) Heuristic-based parsers: These log parsers [4, 7, 11, 18, 28, 16] encode expert domain knowledge into general and effective heuristic rules to identify log templates. Semantic-based log parsers are a recently emerging class of parsers and can achieve higher parsing accuracy by mining semantics from log messages. Specifically, Semantic-based log parsers can be further subdivided into two categories. (1) tuning-based parsers: These log parsers [17, 15, 14, 27, 29] typically need labeled log data for model training or tuning. For example, Uniparser [17] formulates log parsing as a token classification problem, employing Bi-LSTM models for training. VALB [15] formulates log parsing as a sequence tagging problem and uses Bi-LSTM-CRF models for training. (2) prompt-based parsers: These log parsers [25, 12, 9] utilize a designed prompt to query language language models (LLMs) for log parsing. Unlike syntax-based log parsers, LogPTR doesn’t require any handcrafted features or heuristic rules, which can adapt to new log datasets with low effort. Furthermore, compared to semantic-based log parsers, LogPTR can not only identify log templates and variables but also identify the categories of variables, which can aid log understanding and analysis.
@@ 6 CONCLUSION
This paper proposes LogPTR, a simple yet effective variable-aware log parser that can extract static and dynamic parts in logs and further identify categories of variables. The key of LogPTR is formulating log parsing as a text summarization problem and using a pointer mechanism to copy words from the log message and label tokens indicating categories of variables. Experimental results show that LogPTR outperforms the state-of-the-art log parsers on both general log parsing and variable-aware log parsing.